Student are given a list of nine books and told that they will be examined on the contents of five of them.How many combinations of five book are

Answer by mathmate(429)

Question:
Student are given a list of nine books and told that they will be examined on the contents of five of them.How many combinations of five book are

Solution:
Imagine the teacher picks 5 books at random.
Her first book would be one out of 9,
the second book would be one out of 8,
...
the fifth book would be one out of 5, leaving four behind.
Therefore the number of permutations would be 9*8*7*6*5=15120

However, she could have chosen the same five books in many different orders, namely 5*4*3*2*1=120

Therefore, the total number of ways of picking 5 books out of nine, where order does not count, is 15120/120=126.

In general, the number of combinations of r objects taken from n objects is C(n,r)=n!/(r!(n-r)!)=9!/(5!4!)=126, same as above.